Calculate the work (in kJ) when substance initially at 99 kPa, 16 m3 is expanded against a linear spring to 208 kPa, 70 m3.
Answer #1

For linear spring;

In other words, pressure is proportional to the volume.

Average pressure during the process is given as;

Hence, work done is given as;

...(Answer)
Answer #2

To calculate the work done when a substance initially at 99 kPa and 16 m³ is expanded against a linear spring to 208 kPa and 70 m³, we can follow these steps:

Given:

  • Initial pressure, P1=99kPaP_1 = 99 \, \text{kPa}

  • Initial volume, V1=16m3V_1 = 16 \, \text{m}^3

  • Final pressure, P2=208kPaP_2 = 208 \, \text{kPa}

  • Final volume, V2=70m3V_2 = 70 \, \text{m}^3

Assumption:

  • The pressure-volume relationship is linear due to the linear spring, meaning pressure varies linearly with volume.

Calculation:

  1. Determine the linear relationship between pressure and volume:

    Since the relationship is linear, we can express pressure as:

    P=mV+cP = mV + c

    where mm is the slope and cc is the intercept.

    Using the initial and final conditions:

    P1=mV1+c    99=m×16+c(Equation 1)P_1 = mV_1 + c \implies 99 = m \times 16 + c \quad \text{(Equation 1)}P2=mV2+c    208=m×70+c(Equation 2)P_2 = mV_2 + c \implies 208 = m \times 70 + c \quad \text{(Equation 2)}

    Subtract Equation 1 from Equation 2 to solve for mm:

    20899=m×(7016)208 - 99 = m \times (70 - 16)109=m×54109 = m \times 54m=109542.0185kPa/m3m = \frac{109}{54} \approx 2.0185 \, \text{kPa/m}^3

    Substitute mm back into Equation 1 to find cc:

    99=2.0185×16+c99 = 2.0185 \times 16 + c99=32.296+c99 = 32.296 + cc=9932.296c = 99 - 32.296c66.704kPac \approx 66.704 \, \text{kPa}

    Therefore, the pressure-volume relationship is:

    P=2.0185V+66.704P = 2.0185V + 66.704

  2. Calculate the work done:

    The work done during expansion is the area under the PP-VV curve, which can be calculated as:

    W=V1V2PdVW = \int_{V_1}^{V_2} P \, dV

    Substitute the expression for PP:

    W=1670(2.0185V+66.704)dVW = \int_{16}^{70} (2.0185V + 66.704) \, dV

    Evaluate the integral:

    W=[2.0185V22+66.704V]1670W = \left[ 2.0185 \frac{V^2}{2} + 66.704V \right]_{16}^{70}

    Calculate the definite integral:

    W=(2.0185×7022+66.704×70)(2.0185×1622+66.704×16)W = \left( 2.0185 \times \frac{70^2}{2} + 66.704 \times 70 \right) - \left( 2.0185 \times \frac{16^2}{2} + 66.704 \times 16 \right)W=(2.0185×2450+4669.28)(2.0185×128+1067.264)W = \left( 2.0185 \times 2450 + 4669.28 \right) - \left( 2.0185 \times 128 + 1067.264 \right)W=(4945.825+4669.28)(258.368+1067.264)W = \left( 4945.825 + 4669.28 \right) - \left( 258.368 + 1067.264 \right)W=9615.1051325.632W = 9615.105 - 1325.632W=8289.473kPam3W = 8289.473 \, \text{kPa} \cdot \text{m}^3

    Since 1kPam3=1kJ1 \, \text{kPa} \cdot \text{m}^3 = 1 \, \text{kJ}:

    W8289.473kJW \approx 8289.473 \, \text{kJ}


answered by: Shital Garg